February 8, 2017 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, February 8, 2017
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or bjn.vc
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2016/ .

Calibration Tasks

  1. Spring 2016 Recon Checklist
  2. Mcsmear updates
  3. Data Validation
  4. GlueX Detector Performance Document

Upcoming Calibrations

  • Add timing information to FCAL clustering/inner ring calibrations? (Adesh/Matt S.)
  • BCAL TDC calibrations (Andrei)
  • TAGH alignment (Nathan)
  • Tagger energy map (Richard J.?)
  • TOF alignment (Beni/Simon)


  1. Announcements
    • CCDB hourly notification list
    • Adding constants from gluon machines
    • ROOT6 bugs
  2. Collaboration Meeting
  3. Calibration Updates
  4. Simulations
    • sim1.2 progress
  5. AOB


Attending: Sean (NU); Alex A., Adesh, Nathan, Brad, Elton, Simon, Mark I., Beni, Alex B. (JLab); Zisis (Regina); Matt S. (IU); Will M., Naomi, Curtis (CMU); George (Athens); Justin (W&M)

  1. Announcements
    • In addition to the daily email of CCDB updates, there is now a similar hourly email for those who are coordinating changes or want their email inbox filled faster.
    • Beniran into a problem in the past week where he tried to add constants to the CCDB from the gluon machines and this silently failed. He suggested checking the CCDB logs after committing constants to make sure that they get in there. Mark I. will try to reproduce this bug and file a GitHub issue.
    • Sean mentioned that a recent update to the standard ROOT6 version a couple of months ago has caused more monitoring/calibration script errors due to stricter interpreter checking and bugs with TROOT::Reset().
  2. Last call for Collaboration Meeting talks. Mark I. confirmed that Richard will give a talk.
  3. Calibration Updates
    • Sean gave an Overview of the current status of this run. Rough timing calibrations are in which are enough to start looking at final states, although there are potentially some phi-dependent effects. Convergence has been slow due to issues with start counter calibrations. We looked through some of the standard monitoring plots, and noted some features. Further progress requires SC updates and full run statistics.
    • Start Counter - Mahmoud has been working on updating the propagation time calibrations from the previous run. He had run into some problems due to the 2ns beam spacing and paddle-dependent timing shifts. Sean pointed out some issues that could be due to timewalk corrections that need to be updated. The hope is that the current data will be a better source for these calibrations. As Mahmoud is on shift, Sean will work on obtaining some preliminary new timewalk/propogation time corrections today to prepare for large-scale calibration launches.
    • TOF - Beni reported that yesterday he was able to get some updated calibrations in based on run 30218. The overall resolution is ~110 ps, and there is hope for improvement towards getting back to under 100 ps. Sean asked about current problems with matching efficiency on the narrow paddles, Beni mentioned that not all of these calibrations converged and he is looking into it.
    • TAGM - Alex B. reworked his calibration scheme, and discussed the wiki page describing them. This takes multiple passes and needs only ~200k events for a good calibration. The high-level plugin will continue to align the TAGM with the main detector. He added new calibrations to the CCDB
    • TAGH - Nathan found a bug in the TAGH hit merging that caused ~10% of his TPOL analysis jobs to hang. He plans to fix it today, but in the meantime, it can be worked around by disabling the hit merging using the JANA option -PTAGHHit:MERGE_DOUBLES=0. Nathan and Sean noticed some rate dependence in the low-energy counters similar to last Spring, which will require a run-by-run calibration.
    • BCAL - Progress is being made on several fronts. Mark D. noticed that the z positions were miscalibrated, which led to looking at recalibrating the table that calibrates the z positions. George showed some fresh results that there are some systematic shifts that are not understood. These need to be checked with the current data, but the plan is to recalibrate based on these results, and then update the effective velocities and attenuation lengths. The systematics behind these changes still need to be studied. Sean is working on getting these calibrations into the standard procedures. The differences are curious since they should just depend on the difference in hit times between the two ends. Elton had some additional questions that he would ask offline.
    • FCAL - Adesh looked at the pi0 position and width and they looked consistent with what was seen last run. After skims are available, he can look at gain and timing calibrations. He asked Sean to put RF information into the skims for the timing, Sean had done so and will send Adesh information needed to test this. He has confirmed that the LED pulser is running at 10 Hz, and would like the DAQ to run even when there is no beam.
    • Tracking - No news.
    • Train Processing Status - Sean will start a preliminary skim launch after SC and BCAL calibrations are updated, hopefully tonight.
  4. Simulations
    • Mark I. reported that we are 10% done of the 30% of spring 2016 sample size that we planned to simulate. This is enough for people to start looking at. The plan is to reduce the priority of these jobs, in order to prioritize calibration and analysis of the data from this run.
    • Sean pointed out that Thomas B. is beginning to attach the problem of mix-in backgrounds.
  5. AOB